The management of Firebolt Industries Inc. manufactures gasoline and diesel engines through two production departments, Fabrication and Assembly. Management needs accurate product cost information in order to guide product strategy. Presently, the company uses a single plantwide factory overhead rate for allocating factory overhead to the two products. However, management is considering the multiple production department factory overhead rate method. The following factory overhead was budgeted for Firebolt: 1 Fabrication Department factory overhead $614,800.00 2 Assembly Department factory overhead 246,750.00 3 Total $861,550.00 Direct labor hours were estimated as follows: Fabrication Department 5,300 hours Assembly Department 5,250 Total 10,550 hours In addition, the direct labor hours (dlh) used to produce a unit of each product in each department were determined from engineering records, as follows: Production Departments Gasoline Engine Diesel Engine Fabrication Department 2.9 dlh 1.8 dlh Assembly Department 1.8 2.9 Direct labor hours per unit 4.7 dlh 4.7 dlh Required: a. Determine the per-unit factory overhead allocated to the gasoline and diesel engines under the single plantwide factory overhead rate method, using direct labor hours as the activity base.* b. Determine the per-unit factory overhead allocated to the gasoline and diesel engines under the multiple production department factory overhead rate method, using direct labor hours as the activity base for each department.* c. (1) Recommend to management a product costing approach, based on your analyses in (a) and (b). (2) Give a reason for your answer. *If required, round all per-unit answers to the nearest cent. a. Determine the per-unit factory overhead allocated to the gasoline and diesel engines under the single plantwide factory overhead rate method, using direct labor hours as the activity base . If required, round all per-direct labor hours and per-unit answers to the nearest cent. Gasoline engine $ per unit Diesel engine $ per unit b. Determine the per-unit factory overhead allocated to the gasoline and diesel engines under the multiple production department factory overhead rate method, using direct labor hours as the activity base for each department. If required, round all per-unit answers to the nearest cent. Gasoline engine $ per unit Diesel engine $ per unit
Answer #1

Requirement a

Total factory overhead (A) = $ 861,550

Total direct labor hours (B) = 10,550

Single Plantwide overhead rate = A/B = $ 81.66

Particulars

Gasoline Engine

Diesel Engine

Single Plantwide Overhead rate per direct labor hour (A)

$ 81.66

$ 81.66

No. of direct labor hours per unit (B)

4.7

4.7

Factory Overhead allocated per unit (A *B)

$ 383.80

$ 383.80

Requirement b

Particulars

Fabricating Department

Assembly Department

Factory Overhead (A)

$ 6,14,800.00

$ 2,46,750.00

No. of direct labor hours (B)

5300

5250

Overhead rate per direct labor hour (C = A/B)

$ 116.00

$ 47.00

Gasoline Engine

Diesel Engine

Fabricating Department

Direct labor Hours per unit (A)

2.9

1.8

Overhead rate per direct labor hour (B)

$ 116.00

$ 116.00

Overhead allocated per unit (C = A* B)

$ 336.40

$ 208.80

Assembly Department

Direct labor Hours per unit (D)

1.8

2.9

Overhead rate per direct labor hour ( E)

$ 47.00

$ 47.00

Overhead allocated per unit (F = D *E)

$ 84.60

$ 136.30

Total Overhead allocated per unit (C + F)

$ 421.00

$ 345.10

Requirement c

Product costing approach should be applied in multiple department using multiple overhead rates as in requirement b

Even products use same direct labor hours, each of them use different labor hours in various department. Each department should therefore have its own overhead allocation rate. Hence, single plantwide overhead rate is not recommended rather multiple department overhead rate is recommended.

kindly upvote